A juggler tosses three balls alternately vertically upward. Each ball has an initial velocity of 5m/s. How high does each ball rise? How long does each ball remain in the air? Show your solution


1
Expert's answer
2021-05-20T10:08:29-0400

"h=v_0^2/(2g)=5^2/(2\cdot9.8)=1.276\ (m)" . Answer


"t=2t_0"


"v=v_0-gt\to t=(v_0-v)/g=(5-0)/9.8=0.51\ (s)"


"t=2\cdot0.51=1.02\ (s)" . Answer